# Critical Evaluation of Studies Alleging Evidence for Technosignatures in the POSS1-E Photographic Plates

### v1

@@ Acknowledgments
WW received a copy of $V$ in July 2025 while collaborating on a manuscript that was published as Villarroel et al. (2025a) in late September 2025. When voluntarily withdrawing from this paper in July, WW raised early concerns regarding the spatial distribution of SPFs with the lead author (BV). The authors of Villarroel et al. (2025a) stated that “Data will be shared on reasonable request to the corresponding author,” presumably in reference to $S$ or $V$. In the late fall of 2025, coauthors KK and LD (of this study) independently requested copies of these datasets, and were told they are not yet available. Because we have not obtained permission to publish analyses of $V$, we have used and reported here only on the publicly available datasets $R$ and $W$. We are grateful to Doug Buettner, Tejin Cai, Mike Cifone, Eric Keto, Matthew Szenher, and Matthew Szydagis for helpful comments.
@@ Appendix A Ambiguities in dataset definitions
As previously mentioned, the origins and reported sizes of $V$ (used in Bruehl and Villarroel (2025)) and its subsets $V^{\prime}$ and $V^{\prime\prime}$, used for key calculations in Villarroel et al. (2025a), are inconsistent or conflated in Villarroel et al. (2025a). We have listed some of these inconsistent statements here:
  • From Villarroel et al. (2025a), pg 15:“We use the transient candidates from Solano et al. (2022), but with the additional requirement that they have no counterparts within $5^{\prime\prime}$ in Gaia, Pan-STARRS and NeoWise.” The initial candidate selection pool described in Solano et al. (2022), which we have called $S$, already did not included SPFs near Gaia or Pan-STARRS objects (Solano et al., 2022). In any case, removing NeoWISE from $S$ is consistent with our definition of set $P$ in Table 1 (i.e., SWS-W or $A-(a_{0}\cup W)$), which has an expected size of $N=126,412$.

### v2

@@ Acknowledgments
WW received a copy of $V$ in July 2025 while collaborating on a manuscript that was published as Villarroel et al. (2025a) in late September 2025. When voluntarily withdrawing from this paper in July, WW raised early concerns regarding the spatial distribution of SPFs with the lead author (BV). Villarroel et al. (2025a) states that “Data will be shared on reasonable request to the corresponding author,” presumably in reference to $S$ or $V$. In the late fall of 2025, coauthors KK and LD (of this study) independently requested copies of these datasets, and were told they are not yet available. Because we have not obtained permission to publish analyses of $V$, we have used and reported here only on the publicly available datasets $R$ and $W$. KK thanks Tony Gorman for his support of University at Albany Project X (UAPx). We are grateful to Doug Buettner, Tejin Cai, Mike Cifone, Nigel Hambly, Eric Keto, Matthew Szenher, and Matthew Szydagis for helpful comments.
Sets $R$ and $W$ are based on data from the SVO archive of vanishing objects in POSS I red images. The archive was developed in the framework of the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. The system is maintained by the Data Archive Unit of the CAB (CSIC -INTA).
@@ Appendix A Ambiguities in dataset definitions
As previously mentioned, the origins and reported sizes of $V$ (used in Bruehl and Villarroel (2025)) and its subsets $V^{\prime}$ and $V^{\prime\prime}$, used for key calculations in Villarroel et al. (2025a), are inconsistent in Villarroel et al. (2025a). We have listed some of these inconsistent statements here:
  • From Villarroel et al. (2025a), pg 15:“We use the transient candidates from Solano et al. (2022), but with the additional requirement that they have no counterparts within $5^{\prime\prime}$ in Gaia, Pan-STARRS and NeoWise.” The initial candidate selection pool described in Solano et al. (2022), which we have called $S$, already did not include SPFs near Gaia or Pan-STARRS objects (Solano et al., 2022). In any case, removing NeoWISE from $S$ is consistent with our definition of set $P$ in Table 1 (i.e., SWS-W or $A-(a_{0}\cup W)$), which has an expected size of $N=126,412$.